OpenStudy (anonymous):

6) Jodi is studying plant growth rates for her science project. For her project, she selected three bean plants of equal height. Then, for the next five days, she measured the height, in centimeters , of each plant and plotted the values: (1,1)(1,1.8)(1,2.6)(2,2)(2,3.2)(2,4.1)(3,2.8)(3,3.5)(3,5.1)(4,4.1)(4,4.9)(4,5.9)(5,5.8)(5,7)(5,7.2) She drew a line of best fit passing through (1,1)(5,7) on the graph to show one way of calculating the mean growth rate of the plants. What is the slope of the line she drew?

whitemonsterbunny17 (whitemonsterbunny17):

I think this belongs in the Mathematics section. In order to solve this, though, you will need to create a graph, and create a line using those points.

OpenStudy (bioepic):

Slope = (y2-y1)/(x2-x1) Slope = (7-1)/(5-1) = 6/4 = 3/2. 3/2